OpenStudy (anonymous):

55.27 is the same as 55 ____' 12".

OpenStudy (whpalmer4):

0.27 * 60 = take the integer part only to get the minutes, then multiply the decimal part by 60 again to get the seconds (which will turn out to be 12)

OpenStudy (anonymous):

so the missing answer is 16 ? @whpalmer4

OpenStudy (whpalmer4):

Yes. 0.27*60 = 16.2, so 16' 0.2*60 = 12, so 12 "